# Modeling Issues with Eye Tracking Data

### v3

@@ 4 Modeling
A related model to AR1 is a first-order moving average model MA1. When binary data is composed of runs (streaks of 0s and 1s), it suggests that once a saccade occurs (e.g., a state switch), its influence is immediate but might not persist indefinitely across the entire time series. This seems more natural to a eye-tracking time series than an autogressive process. This model can also be fit with GEE in the forms of an $R$ matrix having a tri-diagonal structure, also known as a banded Toeplitz matrix of width 1. For this correlation matrix, only the off-diagonals immediately above and below the main diagonal have a nonzero value of $\phi_{M}$. The estimating equations are identical to those of the AR1 model.
@@ 6 Survival Modeling
RLE modeling of eye tracking responses is designed to model the 0-1 category frequencies, rather than the 0-1 data individually. This strategy was suggested previously in the discussion of Table 1. For this purpose, binomial modeling can be used in the framework of a GLMM. However, it is convenient to implement this analysis relative to runs within a subject-item cluster. In fact, the runs are treated as discrete instances of gaze that do not correspond to a linear process, either manifest or latent.
For RLE compression, the blocks correspond to $Y=0$ or to $Y=1$. After compression, each episode can be represented by an interval $Time_{\text{start}},Time_{\text{stop}}]$ and an associated event indicator that marks whether a transition to the opposite gaze state occurred at the end of the interval. This representation lends itself to an alternating two–state survival framework, where separate transition hazards are modeled for leaving the target $1\rightarrow 0$ and entering the target $0\rightarrow 1$ [kleinbaum2012survival, therneau2024survival]. Because RLE compression preserves the dwell structure, the resulting episode durations provide suitable survival intervals for Cox regression. The COX model is:
@@
In the fitted model, the baseline hazard is stratified by transition type so that the temporal risk profiles for leaving and entering the target object may differ freely. Predictors such as Privileged and t are allowed to influence each transition differently through interaction terms with the transition indicator. This structure has interpretable effects: the main coefficients describe how covariates affect the hazard of leaving the target, while the interaction coefficients quantify how those effects change during target reacquisition. Main and transition effects and standard errors can be combined to represent total effects. For this analysis, the main focus is on the total effect for the $0\rightarrow 1$ transition. The survival model with RLE data provides a computationally efficient modeling alternative to AR1 with uncompressed data.
The model incorporates a Cox proportional hazards framework, the process being modeled is not survival in the conventional sense. The data consist of a binary time series in which each subject–item trajectory alternates repeatedly between two states, with each interval representing the duration spent in a given state before switching to the other. In this context, the COX model is used as a flexible estimator of state-specific switching intensities rather than time to a terminal event. By stratifying on transition direction, separate baseline hazards arfe used for $0\rightarrow 1$ and $1\rightarrow 0$ transitions, so that covariate effects are interpreted as changes in the instantaneous rate of leaving the current state.
@@ 7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmgee [Vanegas2023] was used for analysis of the AR1/MA1 models, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 and MA1 models produced positive definite $R$ matrices; however, their determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ and a ridge of 1e-5 was applied to $R$. In both cases, estimates were nearly identical to those produced in runs where $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was missing by fiat for the last run. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, MA1, and COX models.
@@
Privileged 0.077 0.052 0.070 0.083 0.053
Contrast -0.079 -0.057 -0.068 -0.013 -0.046
Time 0.854 0.813 0.831 0.498
Priv*Time -0.012 0.001 0.006 0.012
Contrast*Time -0.015 -0.005 -0.012 0.028
@@ 7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and the transition interaction ($y=0\rightarrow y=1)$, and likewise for $Contrast$. It can be seen that the GLM and MA1 methods provide similar estimates of experimental effects, and the AR1 and COX methods also have similar estimates. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model, and the $Time$ coefficient is strikingly dissimilar to those of the other models. The GLM, AR1, and MA1 models indicate a strong positive relationship between time and the likelihood of being on-target. The variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. For this reason, the COX method did not produce intercepts or time interactions.
It is interesting that the AR1 and MA1 models resulted in moderately different estimates of experimental effects. In this regard, the MA1 model is most similar to the GLM model. Recall that the AR1, MA1, and COX approaches employ variance modeling alternatives to error correlation, while no adjustment for autocorrelation is present in the GLM model. Both models also indicate negligible interactions over time for the experimental variables.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM provides the lowest, and the AR1/MA1, COX models provide the highest for the experimental valriables. The LAG model produces SEs about double of those for the GLM model, but even so, LAG SEs are half the size those for AR1, MA1, and COX. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator. Still, the SEs for the AR1, MA1, and COX models are very similar. This may imply that these three models implicitly weight transition nieghborhoods higher than other areas of the time series.
@@
Privileged 0.007 0.034 0.037 0.014 0.030
Contrast 0.004 0.021 0.023 0.009 0.018
Time 0.003 0.014 0.015 0.006
Priv*Time 0.007 0.033 0.035 0.015
Contrast*Time 0.005 0.020 0.022 0.009
@@ 7.3 Random Effects and $\phi$
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, MA1, and COX models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were similar. For the AR1 and MA1 model, the estimated residual correlations are extremely high. However, this is not unexpected given the median run length of 35 (out of 112). This level of $\phi$ for AR1 and MA1 raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient.
@@
MA1 0.95 (0.963) 1.001
  • Note. Ridge estimates are reported for AR1 and MA1 with $\phi=.95$. Estimated values of $\phi$ are given in parentheses.
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, AR1/MA1 and COX provide more conservative alternatives for estimating fixed effects. The results of the AR1 and COX models raise important and unresolved questions. First, the finding that these two approaches provide similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions.
@@
Of course, the similarity of the AR1 and COX models does not imply they are correct. Rather it raises the question of how eye-tracking time series can be modeled. The AR1 and MA1 models are impractical for longer time series because the determinant of $R$ approaches zero. While a upper boundary can be set manually, this is purely an ad hoc solution. The survival model, or similar models, may provide a way to avoid near singularities due to serial correlation. Moreover, these models are not truly time series models but rather error-correlation models. Based on results from the current study, it is not clear that time series models are require or even suitable for eye-tracking data.
@@
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1/MA1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient seems to have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction the subject variance component as shown in Table 4.
@@
The results of this study collectively support the existence of experimental effects for the variable $Contrast$. For the AR1, MA1, and COX models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1/MA1 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength, this demonstrates model assumptions can make substantive differences in both effect size and statistical significance.

### v4

@@ 4 Modeling
A related model to AR1 is a first-order moving average model (MA). When binary data is composed of runs (streaks of 0s and 1s), it suggests that once a saccade occurs (e.g., a state switch), its influence is immediate but might not persist indefinitely across the entire time series. This seems more natural to a eye-tracking time series than an autogressive process. This model can also be fit with GEE in the forms of an $R$ matrix having a banded Toeplitz structure. I chose a width of 25 which is approximately the median band length for subjects with at least four runs. For this correlation matrix, $\phi_{M}=.95$. The estimating equations are identical to those of the AR1 model. I use this asa point of comparison only. Subjects’ runs vary in number and length, so setting a fixed bandwidth is highly arbitrary.
@@ 6 Survival Modeling
RLE modeling of eye tracking responses could be used to model the 0-1 category frequencies, rather than the 0-1 data individually. This strategy was suggested previously in the discussion of Table 1. While binomial modeling can be used in the framework of a GLMM. However, it this results in a serial correlation of $\rho=-1$. It is more interesting to treat the runs treated as discrete gaze episodes that occur over time. For RLE compression, the blocks correspond to $Y=0$ or to $Y=1$. After compression, each episode can be represented by an interval $Time_{\text{start}},Time_{\text{stop}}]$ and an associated event indicator that marks whether a transition to the opposite gaze state occurred at the end of the interval. This representation lends itself to an alternating two–state survival framework, where separate transition hazards are modeled for leaving the target $1\rightarrow 0$ and entering the target $0\rightarrow 1$ [kleinbaum2012survival, therneau2024survival]. Because RLE compression preserves the dwell structure, the resulting episode durations provide suitable survival intervals for Cox regression. The COX model is:
@@
In the fitted model, the baseline hazard is stratified by transition type so that the temporal risk profiles for leaving and entering the target object may differ freely. Predictors such as Privileged and t are allowed to influence each transition differently through interaction terms with the transition indicator. This structure has interpretable effects: the main coefficients describe how covariates affect the hazard of leaving the target, while the interaction coefficients quantify how those effects change during target reacquisition. Main and transition effects and standard errors can be combined to represent total effects. For this analysis, the main focus is on the total effect for the $0\rightarrow 1$ transition. The survival model with RLE-encoded data provides a computationally efficient modeling alternative to AR1 with uncompressed data.
The model incorporates a Cox proportional hazards framework, the process being modeled is not survival in the conventional sense. The data consist of a binary time series in which each subject–item trajectory alternates repeatedly between two states, with each interval representing the duration spent in a given state before switching to the other. In this context, the COX model is used as a flexible estimator of state-specific switching intensities rather than time to a terminal event. By stratifying on transition direction, separate baseline hazards are used for $0\rightarrow 1$ and $1\rightarrow 0$ transitions, so that covariate effects are interpreted as changes in the instantaneous rate of leaving the current state.
@@ 7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmtoolbox [Vanegas2023] was used for analysis of the AR1/MA25 models, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 model produced positive definite $R$ matrices; however, the determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ and a ridge of 1e-5 (or 1e-2 for MA25) was applied to $R$. For AR1, estimates were nearly identical to those produced in runs where $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was missing by fiat for the last run. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, MA25, and COX models.
@@
Privileged 0.077 0.052 0.052 0.083 0.053
Contrast -0.079 -0.057 -0.041 -0.013 -0.046
Time 0.854 0.813 0.736 0.498
Priv*Time -0.012 0.001 -0.016 0.012
Contrast*Time -0.015 -0.005 0.007 0.028
@@ 7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and the transition interaction ($y=0\rightarrow y=1)$, and likewise for $Contrast$. It can be seen that the GLM and AR1/MA25 methods provide similar estimates of experimental effects, and the AR1 and COX methods also have similar estimates. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model, and the $Time$ coefficient is strikingly dissimilar to those of the other models. The GLM, AR1, and MA25 models indicate a strong positive relationship between time and the likelihood of being on-target. The variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. For this reason, the COX method did not produce intercepts or time interactions.
It is interesting that the AR1/MA25 models resulted in similar estimates of experimental effects. This is consistent with the the utility of GEE esimtation providing robust estimates given misspecification of the error correlation. Recall that the AR1, MA25, and COX approaches employ variance modeling alternatives to error correlation, while no adjustment for autocorrelation is present in the GLM model.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM provides the lowest, and the AR1/MA25, COX models provide the highest for the experimental variables. The LAG model produces SEs about double of those for the GLM model, but even so, LAG SEs are half the size those for AR1/MA25 and COX. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator. Still, the SEs for the AR1/MA25 and COX models are similar.
@@
Privileged 0.007 0.034 0.036 0.014 0.030
Contrast 0.004 0.021 0.021 0.009 0.018
Time 0.003 0.014 0.027 0.006
Priv*Time 0.007 0.033 0.066 0.015
Contrast*Time 0.005 0.020 0.038 0.009
@@ 7.3 Random Effects and $\phi$
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, MA25, and COX models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were similar. For the AR1, the estimated residual correlation was extremely high. However, this is not unexpected given the median run length of 35 (out of 112) across all subjects. This level of $\phi$ raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient.
@@
MA25 0.95 (—) 0.967
  • Note. Ridge estimates are reported for AR1 and MA1 with $\phi=.95$. Estimated values of $\phi$ are given in parentheses, but $\phi$ as not estimated for MA25.
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, the AR1/MA25 and COX provide more conservative alternatives for estimating fixed effects. This result raises important and unresolved questions. First, the finding that these two approaches provide similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions.
@@
Of course, the similarity of the AR1/MA25 and COX models does not imply they are correct. Rather it raises the question of how eye-tracking time series can be modeled. The AR1 and moving average models are impractical for longer time series because the determinant of $R$ approaches zero. While a upper boundary can be set manually, this is purely an ad hoc solution. The survival model, or similar models, may provide a way to avoid near singularities due to serial correlation. Moreover, these models are not truly time series models but rather error-correlation models. Based on results from the current study, it is not clear that time series or error correlztion models are required or even suitable for eye-tracking data.
@@
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient seems to have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction the subject variance component as shown in Table 4.
@@
The results of this study collectively support the existence of experimental effects for the variable $Contrast$. For the AR1, MA25, and COX models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1/MA25 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength, this demonstrates model assumptions can make substantively important differences.